# Jing Liu EnQi Lian

### v1

@@ 4.1.2 Potential of Initial Latent
In the open-source audio-to-audio method, we encode the original audio through a VAE and incorporate it into randomly initialized latent representations. To ensure fair comparison, we maintain consistency across all methods using Stable Audio Open, including audio-to-audio, generating from scratch, and our inversion model. Since our objective is to examine the potential of initial latent representations, we deliberately avoid introducing additional adaptive editing schemes during the sampling process. Specifically, we do not intervene in the inverted latent representation throughout sampling but maintain methodological consistency by solely modifying the initial latent assignment. For evaluation, we employ common knowledge prompts, those with sufficient training data during pre-training, as text inputs. We apply identical text prompts and original audio (except for generating from scratch) across all comparative methods, ensuring that the same sampler and parameters are applied to maintain experimental validity.
Stable Audio 2 is a commercial version, and both its built-in audio-to-audio function and the sampling process remain black boxes to us. While it claims to have incorporated this functionality on top of the previous version (Evans et al., 2024a), it is unclear whether it employs the same methodology as the open-source approach. Additionally, this version demonstrates improved model capabilities compared to the previous version, so we include it as an additional reference point.
There are subtle differences among the generated audio from the open-source model, and it is practically impossible to distinguish the original sound components across all five samples (see Figure 3). As expected, audio generated by other methods should exhibit greater similarity to the original sound than samples generated from scratch; however, all outputs of Stable Audio Open sound similar regardless of the original audio content. Though the sounds generated by the open-source and commercial versions vary, both adhere to the text prompt ”An airplane takes off.”
As discussed in the methodology section, this phenomenon arises due to the noise control coefficient and signal strength coefficient, which make the noise guided by text prompts dominate generation during the early sampling stages. This is particularly evident in generation from scratch, where the influence of the initial latent representation is relatively weak. These findings indicate that additional controls over the sampling trajectory need to be integrated into the sampling process to achieve ideal reconstruction results.
4.1.3 Latent Intervention Across Sampler Orders
Figure A.2 demonstrates reconstructed samples when intervening with inverted latent codes at timesteps 0, 80, 120, 140, and 180, compared to the sample $x_{0}$ in the last column. The three rows employ first-, second-, and third-order samplers respectively. First-order sampling is equivalent to DDIM. It is evident that features illustrated in the third-order reconstruction are relatively vague compared to the other two. Although we insert the latent at certain timesteps, higher-order methods compute with interpolations, therefore, the influence of these interpolations persists. Consequently, the intervened latents do not dominate the reconstruction to the same extent as in first- and second-order sampling. At the end of sampling, sigma, the noise scheduler, becomes very small, and the predicted noise cannot compete with the inserted latents that preserve most of the features. This indicates that initially, the prompt controls the generation process, where sigma is fairly large.
@@ 4.2 Reference Audio Reconstruction
The audio reconstructed through model fine-tuning restores the reference sound fairly completely, but some loss and confusion of details still exist. As shown in the spectrograms of Figure 4, the reconstruction process generates a layer of blurry background noise, with the sound’s stylistic features enhanced over this background noise. This phenomenon is most evident in the low-frequency portions. As we analyzed earlier, the VAE is designed to prioritize audibility and tends to smooth the regions around sampling points, leading to such artifacts in the sound. While the sound lacks clarity in details, it is acceptable overall. Therefore, we believe that the noise prediction module has acquired the capability to reconstruct text descriptions into audio with reference sound characteristics.
Figure 4: Spectrograms of (a) reference audio and (b) reconstructed reference audio.
4.3 Audio Fusion
We observe that inverted latent representation intervention at timesteps within a certain interval produces audio with optimal fusion results, indicating the existence of a fusion window. Within this window, as the intervention timestep decreases, the fused audio gradually transitions from resembling the reference audio to resembling the original audio in terms of auditory characteristics. Specifically, fusion represents a process of mutual balance between the reference sound and the original sound, where identifying the equilibrium point between the two within the window becomes crucial. In this section, we present the results of audio fusion and subjective evaluation of the fusion effectiveness.
(a) Set 1.
(b) Set 2.
(c) Set 3.
(d) Set 4.
(e) Set 5.
Figure 6: Sound fusion results. The first, second, and third columns represent the original sound, reference sound, and fused sound, respectively.
4.3.1 Results
We used text descriptions as control conditions for the noise prediction network, intervening with inverted latents during the sampling process at regular intervals of 10 timesteps, starting from $t=0$. When intervention occurs near the end of sampling, the sound is almost indistinguishable from the original sound, and the reference sound components are difficult to recognize. We therefore intervened with latent representations in the middle-to-late stages of the sampling process to achieve optimal fusion results. Figure 6 shows fusion examples.
4.3.2 Perceptual Evaluation of Fusion Effects
We received a total of 48 valid questionnaires, with most respondents frequently participating in music activities. Professional and vocational practitioners constitute the largest proportion, as shown in Figure 7.
Figure 7: Distribution of respondents’ level of music engagement.
We observed in our experiments that as the intervention timestep decreased, the auditory similarity of the fused sound gradually transitioned from resembling the reference sound to being more similar to the original sound. We selected fusion examples where the intervention timesteps are positioned within the fusion window, and the similarity distribution from subjective evaluation aligns well with our experimental observations. We used the fusion examples from the previous section to query respondents. In Figure 8, ”Audio 1” represents the original audio, ”Audio 2” represents the reference audio, and ”Audio 3” represents the fused audio; the first through fifth audio sets correspond to Set 1 through Set 5 in Figure 6, respectively. Overall, respondents rated the fusion effects favorably, indicating that we successfully achieved the goal of sound fusion.
@@ 5 Conclusion
In this paper, we propose inversion models based on DPMSolver++ with two variations: SDE and ODE. We make the integral term as constant, thereby avoiding the circular dependency of expanding the noise prediction within the sampling formula. Additionally, the deterministic formulation allows for straightforward computation of the noise map at each inversion step.
Our approach retains the flexibility of prompt guidance for both positive and negative guidance. It is computationally efficient. We explain the necessity of removing stochastic terms from the inversion model, demonstrating that approximating inversion with forward diffusion is not applicable for DPMSolver++ sampling phase.
@@
Previous inversion approaches typically computed only the initial latent, and our experiments show that the initial latent alone is insufficient to capture features of the original sample. Therefore, for editing applications, it must be combined with additional control approaches, such as optimized prompt embeddings that control the sampling trajectory, or by intervening in the sampling phase as we propose. In the latter case, we do not necessarily need to invert all the way to the final timestep.
@@
Inversion is employed to reconstruct the auditory features of the original sound, while fine-tuning helps preserve the features of the reference sound through prompting. An optimal interval exists during which intervention achieves effective fusion. As the timestep decreases, the auditory features of the fused sound shift from those of the reference sound to those of the original sound. Higher-order samplers provide a wider fusion interval. Due to the lack of perceptual metrics for general audio, and because our sound fusion approach does not conform to specific physical constraints, as indicated by participant feedback in Appendix B, objective evaluation of our fusion results presents significant challenges. Therefore, we primarily rely on subjective auditory evaluation. The results validate the effectiveness of the sound fusion and inversion methods proposed in this work.
Appendix A Latent Intervention across Sampler Orders: Additional Results
This section provides comprehensive visual results for the latent intervention experiments across different sampler orders discussed in Section 4.1.3. Figure A.2 shows detailed spectrograms comparing first-, second-, and third-order sampling with interventions at various timesteps.
@@ Appendix A Latent Intervention across Sampler Orders: Additional Results
Figure A.2: SDE reconstruction after inversion with null prompt guidance. Columns 1-5: intervention at t=0, 80, 120, 140, 160. Column 6: original audio through VAE.
Appendix B Subjective Analysis on Fusion Results
Fusion is achieved through music generation models, making it difficult to control granularity at the physical or acoustic level. Professional electronic music analysis reveals distinct fusion patterns across the examples.
In the first example, the original audio concentrates in high frequencies, while the reference audio shows evenly distributed frequencies with prominent highs. The fused audio resembles the reference audio’s frequency distribution but adopts the ”melodic direction” of the original audio.
@@ Appendix B Subjective Analysis on Fusion Results
Appendix C Derived Noise Maps
At the end of Section 3.3, we derived the noise map at timestep $t$ given $x_{0}$ and $x_{t}$ in Eq. 9. Below, we demonstrate the relationships between inverted latent interpolations and derived noise maps, as well as the interpolations in the sampling process that utilize the inverted latent and its corresponding noise map at each step:
(c) Restoration of the original sound through sampling using inverted latent interpolations and noise maps (timesteps decreasing from left to right).
Figure C.2: Sampling from inverted latents and derived noise maps, such that the sampling trajectory coincides with the inversion trajectory.
Appendix D VAE Blurs Acoustic Details
Initially, the audio is encoded by the Variational AutoEncoder (VAE) to derive $x_{0}$. Figure D.1 presents a comparative analysis: the left panel depicts the spectrogram of the original audio signal; the center shows the spectrogram of audio reconstructed through VAE encoding-decoding processes; the right shows audio reconstructed from inverted $x_{t}$ using null prompt guidance. While GANs typically produce high-definition samples, VAEs generate comparatively smoother outputs. Constrained by these limitations, VAE encoding-decoding procedures result in perceptible attenuation of acoustic details, manifested in the spectrogram as indistinct boundaries and homogenized frequency band stratification. This phenomenon derives from a fundamental VAE property: latent space continuity. To enhance generation fluidity and minimize artifacts and distortion, proximate points within the VAE latent space decode to similar outputs, though these points may not correspond to actual encoded inputs. The right-side representation, when compared to the center, exhibits additional loss of features in upper frequency bands that possess low energy. Nevertheless, the VAE preserves the macro-structure, with pronounced periodic characteristics remaining intact. The left-center comparison serves to demonstrate inherent model constraints, isolating these from potential effects attributable to the inversion models themselves. Consequently, evaluations in this paper compare reconstructions through VAE encoding-decoding with original samples, rather than comparing with the original samples directly.
Figure D.1: Spectrogram of the original audio (left), spectrogram of audio reconstructed through VAE encoding-decoding (center), and spectrogram of audio reconstructed by sampling from the inverted latent intervened at step 100/ 200 (right).

### v2

@@ 4.1.2 Comparisons with Audio-to-Audio Approaches
For the open-source audio-to-audio method, we encoded the original audio through a VAE and incorporated it into randomly initialized latent representations. To ensure a fair comparison, we evaluated the SDE inversion model independently on the initial latent that deliberately avoided introducing additional adaptive editing schemes during the sampling process. We employed common knowledge prompts—those well-represented in pre-training data—as text inputs for all comparative methods. Stable Audio 2 is a commercial versionstable_audio_2 whose audio-to-audio function and sampling process remain black boxes. Thus, we included it as an additional reference for comparison.
There are subtle differences among the generated audio from the open-source model, and it is practically impossible to distinguish the original audio components across all four samples (see Figure 2). As expected, audio generated by other methods should exhibit greater similarity to the original audio than samples generated from scratch; however, all outputs of Stable Audio Open sounded similar regardless of the original audio content.
Figure 2: Comparison of audio generated from initial latents under different methods. Prompt: "An airplane takes off." (a) Original audio through VAE. (b) Sample from scratch. (c) Our inversion approach. (d) Audio-to-audio adaptation on Stable Audio Open 1.0. (e) Built-in audio-to-audio of Stable Audio 2.0.
This phenomenon arises due to the noise control coefficient and signal strength coefficient, which make the noise guided by text prompts dominate generation during the early sampling stages. This is particularly evident in generation from scratch, where the influence of the initial latent representation is relatively weak. These observations indicate that additional controls over the sampling trajectory need to be integrated into the sampling process to achieve ideal fusion results.
@@ 4.2 Reference Audio Reconstruction
We fine-tuned Stable Audio Open for reference audio reconstruction. While the reconstructed audio restores the reference audio fairly completely, some detail loss and artifacts still persist. As shown in the spectrograms in Figure 3, the reconstruction process generates a layer of blurry background noise, with the audio’s stylistic features enhanced against this background noise. This phenomenon is most evident in the low-frequency regions. The reason for this can be found in Appendix E. While the audio lacks detail clarity, it is acceptable overall. Therefore, we believe that the denoising module has acquired the capability to reconstruct text descriptions into audio with reference audio characteristics.
Figure 3: Spectrograms of reference audio (left) and reconstructed reference audio (right).
4.3 Sound Fusion
We observed that inverted latent representation intervention at timesteps within a certain interval produced audio with optimal fusion results, indicating the existence of a fusion window. Within this window, as the intervention timestep decreases, the fused audio gradually transitions from resembling the reference audio to resembling the original audio in terms of auditory characteristics. Specifically, fusion represents a process of mutual balance between the reference audio and the original audio, where identifying the equilibrium point between the two within the window becomes crucial.
We received a total of 48 valid questionnaires, with most respondents frequently participating in music activities. Music students make up the largest group with 22 participants (45.83%); music professionals and enthusiasts each represent equal portions of the sample, with 12 participants each (25.0%); only 4.17% are infrequent listeners.
We selected fusion examples where the intervention timesteps are positioned within the fusion interval. In Figure 4, "Audio 1" represents the original audio, "Audio 2" represents the reference audio, and "Audio 3" represents the fused audio; the first through fifth audio sets correspond to Set 1 through Set 5 in Figure D.2, respectively. Overall, respondents rated the fusion effects favorably, indicating that we successfully achieved the goal of sound fusion.
@@ 5 Conclusion
In this paper, we propose inversion models based on DPMSolver++ with two variations: SDE and ODE. We treat the integral term as a constant, thereby avoiding the circular dependency (detailed in Appendix A) that arises from expanding the noise prediction within the sampling formula. Additionally, the deterministic formulation enables straightforward computation of the noise map at each inversion step. We explain the necessity of removing stochastic terms from the inversion model, demonstrating that approximating inversion with forward diffusion is not applicable for DPMSolver++ sampling phase.
@@
Our experiments show that the initial latent alone is insufficient to preserve the features of inverted audio against text guidance in sampling. Therefore, it must be combined with additional control mechanisms, such as optimized prompt embeddings that guide the sampling trajectory, or intervention in the sampling phase as we propose for fusion purposes. In the latter approach, we do not necessarily need to invert all the way to the final timestep.
@@
Inversion is employed to reconstruct the auditory features of the original audio, while fine-tuning helps preserve the features of the reference audio through prompting. An optimal interval exists during which intervention achieves effective fusion. As the timestep decreases, the auditory features of the fused audio shift from those of the reference audio to those of the original audio. Higher-order samplers provide a wider fusion interval. Due to the lack of perceptual metrics for general audio, and because our audio fusion approach does not conform to specific physical constraints, as indicated by participant feedback in Appendix D, objective evaluation of our fusion results presents significant challenges. Therefore, we primarily rely on subjective auditory evaluation. The results validate the effectiveness of the sound fusion and inversion methods proposed in this work.
Appendix A Predicted Noise Introduces Circular Dependency
Lu et al. DPMSolver++ proved that many high-order samplers are unstable at large guided scales, which amplified the derivatives of the noise prediction model and affect the convergence range of the ODE solver. The authors proposed replacing the noise prediction term with the model output $x_{\theta}$ to achieve more stable convergence.
Previous research performed inversion on predicted noise. However, expanding the model output with the predicted noise expression results in circular dependency. To address this circular dependency issue intuitively, we substitute the general formula in Eq. 2 into the second-order SDE case and target the latent at the next timestep $r$ as below:
$\boldsymbol{x}_{r}=\frac{2r_{1}(\sigma_{r}^{2}+1)}{\alpha_{t}(1-e^{-2h})}\left(\boldsymbol{x}_{t}-\frac{\sigma_{t}}{\sigma_{s}}e^{-h}\boldsymbol{x}_{s}+\frac{\alpha_{t}(1-2r_{1})(1-e^{-2h})}{2r_{1}}x_{\theta}^{(s)}+\frac{\sigma_{r}\alpha_{t}(1-e^{-2h})}{2r_{1}\sqrt{\sigma_{r}^{2}+1}}\epsilon^{(r)}_{\theta}\right)$
Appendix B Derived Noise Maps
We have inversion models for both SDE and ODE variants of DPMSolver++. We can further derive the noise map at $t$ from Eq. 1:
$\boldsymbol{\epsilon}^{(t)}=\frac{\boldsymbol{x}_{t}-\alpha_{t}\boldsymbol{x}_{0}}{\sigma_{t}}$
We demonstrate the relationships between inverted latent interpolations and derived noise maps, as well as the interpolations in the sampling process that utilize the inverted latent and its corresponding noise map at each step in Figure B.2.
(c) Restoration of the original audio through sampling using inverted latent interpolations and noise maps (timesteps decreasing from left to right).
Figure B.2: Sampling from inverted latents and derived noise maps, such that the sampling trajectory coincides with the inversion trajectory.
Appendix C Latent Intervention Across Sampler Orders
Figure C.2 demonstrates reconstructed samples when intervening with inverted latent codes at timesteps 0, 80, 120, 140, and 180, compared to the sample $\boldsymbol{x}_{0}$ in the last column. The three rows employ first-, second-, and third-order samplers respectively. First-order sampling is equivalent to DDIM. It is evident that features illustrated in the third-order reconstruction are relatively vague compared to the other two. Although we insert the latent at certain timesteps, higher-order methods compute with interpolations, therefore, the influence of these interpolations persists. Consequently, the intervened latents do not dominate the reconstruction to the same extent as in first- and second-order sampling. At the end of sampling, sigma, the noise scheduler, becomes very small, and the predicted noise cannot compete with the inserted latents that preserve most of the features. This indicates that initially, the prompt controls the generation process, where sigma is fairly large.
@@ Appendix C Latent Intervention Across Sampler Orders
Figure C.2: SDE reconstruction after inversion with null prompt guidance. Columns 1-5: intervention at t=0, 80, 120, 140, 160. Column 6: original audio through VAE.
Appendix D Subjective Analysis on Fusion Results
Fusion is achieved through music generation models, making it difficult to control granularity at the physical or acoustic level. This section presents professional electronic music analysis that reveals distinct fusion patterns across the examples.
In the first example of Figure D.2, the original audio concentrates in high frequencies, while the reference audio shows evenly distributed frequencies with prominent highs. The fused audio resembles the reference audio’s frequency distribution but retains the "melodic direction" of the original audio.
@@ Appendix D Subjective Analysis on Fusion Results
(a) Set 1.
(b) Set 2.
(c) Set 3.
(d) Set 4.
(e) Set 5.
Figure D.2: Audio fusion results. The first, second, and third columns represent the original audio, reference audio, and fused audio, respectively.
Appendix E VAE Blurs Acoustic Details
Initially, the audio is encoded by the VAE to derive $x_{0}$. Figure E.1 presents a comparative analysis. While GANs typically produce high-definition samples, VAEs generate comparatively smoother outputs. Constrained by these limitations, VAE encoding-decoding procedures result in perceptible attenuation of acoustic details, manifested in the spectrogram as indistinct boundaries and homogenized frequency band stratification. This phenomenon derives from a fundamental VAE property: latent space continuity. To enhance generation fluidity and minimize artifacts and distortion, proximate points within the VAE latent space decode to similar outputs, though these points may not correspond to actual encoded inputs. The right-side representation, when compared to the center, exhibits additional loss of features in upper frequency bands that possess low energy. Nevertheless, the VAE preserves the macro-structure, with pronounced periodic characteristics remaining intact. The left-center comparison serves to demonstrate inherent model constraints, isolating these from potential effects attributable to the inversion models themselves. Consequently, evaluations in this paper compare reconstructions through VAE encoding-decoding with original samples, rather than comparing with the original samples directly.
Figure E.1: Spectrogram of the original audio (left), spectrogram of audio reconstructed through VAE encoding-decoding (center), and spectrogram of audio reconstructed from inverted $x_{t}$ at step 100/ 200 using null prompt guidance (right).